Let Yi, Y2...., Y4o be a sequence of exponentially distributed random variable with parameter λ=.4, and Y=Σ Y/40, Find P(Y<2) 40

I know this could be solve using the gamma function, is there other way to solve this problem? if so could someone show me how to do it both ways, meaning using the gamma function and any other way possible. please be as detailed as possible. Thank you and I'll rate
Answer #1

Y is sample mean
since n = 40 > 30
we can approximate it using normal distribution

E(Y) = 1/lambda = 1/ 0.4 = 2.5
sd(Y) = 2.5/sqrt(40) = 0.39528

P(Y < 2)
= P(Z <(2 - 2.5)/0.39528)
= P (Z<−1.26)=0.1038